OpenStudy (anonymous):

whats the answer to this problem (8+3i)2

OpenStudy (anonymous):

is the 2 at the end an exponent?

Directrix (directrix):

(8+3i)^2 = (8+3i)*(8+3i) = 64 + 24i + 24i +(3i)^2 = 64 + 48i + 9i^2 = 64 + 48i + 9 (-1) = 64 + 48i - 9 = 55 + 48i.